OpenStudy (anonymous):

Consider the line 5x+2y=−7 . What is the slope of a line parallel to this line? What is the slope of a line perpendicular to this line?

OpenStudy (anonymous):

@triciaal @texaschic101 @ganeshie8 @Walkerman15 @happycam31 @yoitsren

OpenStudy (anonymous):

@dejmccray @eighthourlunch @godinez1243 @undeadknight26 @poopsiedoodle

OpenStudy (anonymous):

Have you solved for y yet?

OpenStudy (anonymous):

i don't know if i'm right but for my i got 2y=-5x-7

OpenStudy (anonymous):

i divide by 2 and got y=-5/2x-3.5

OpenStudy (texaschic101):

you need to put the equation in y = mx + b form. The slope is in the m position. A parallel line wil have the same slope and a perpendicular line will have the negative reciprocal slope.

OpenStudy (anonymous):

i already have a negative slope which is -5/2

OpenStudy (anonymous):

so if it perppendicular the slope is postive

OpenStudy (texaschic101):

5x + 2y = -7 --- subtract 5x from both sides 2y = -5x - 7 -- divide both sides by 2 y = -5/2x - 7/2 your correct...-5/2 is your slope. Therefore, your parallel line will have -5/2 as the slope also. A perpendicular line will have a negative reciprocal slope. All that means is flip " the slope and change the sign. o the perpendicular slope is 2/5 (see how I flipped the slope and changed the sign) any questions ?

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

can you help me find the parallel slope

OpenStudy (texaschic101):

the parallel sloe will be the same slope as in the original equation. Simply because parallel lines have the same slope

OpenStudy (texaschic101):

oops..typo...slope

OpenStudy (anonymous):

is it -5/2

OpenStudy (texaschic101):

parallel slope = -5/2 perpendicular slope = 2/5

OpenStudy (anonymous):

thank you @texaschic101

OpenStudy (texaschic101):

not a problem...glad to help :)